OpenStudy (anonymous):

Help Please!! I will give a medal and become a fan! Lucas opened a savings account 20 years ago with a deposit of $3,815.24. The account has an interest rate of 3.3% compounded quarterly. How much interest has Lucas earned?

OpenStudy (phi):

do you have a formula for compound interest ?

OpenStudy (anonymous):

Yeah and I used it but I still didnt figure it out.

OpenStudy (anonymous):

Nevermind I got it.

OpenStudy (phi):

I would use P = P0 * (1+i/n)^nt P0 is the starting amount 3,815.24 n=4, quarterly means you divide the yearly interest rate by 4 t= 20 years (so nt is 4*20 or 80) i is 3.3% changed to a decimal. 0.033 what did you get ?

OpenStudy (anonymous):

$3,546.50

OpenStudy (phi):

Start with the formula P = P0 * (1+i/n)^nt replace the letters with their values. Don't do any arithmetic, just write in the numbers, to see if you are starting off correctly

OpenStudy (phi):

The interest is (1+i/n)^nt -1 all times P0 you get your number

OpenStudy (anonymous):

\[3,815.24(1+\frac{ 3.3% }{ 4 })^{4\times20}\]

OpenStudy (anonymous):

Oh ok

OpenStudy (phi):

your equation should use 3.3% changed to a decimal. 3.3% is the same as 3.3/100 or 0.033 you should use 0.033 if you do that you will find the total amount, principal plus interest. if you subtract off the principal, you get the interest